# TimeBench: A Comprehensive Evaluation of Temporal Reasoning Abilities in Large Language Models

### v1

@@ 7.1 Temporal Reasoning
Numerous research efforts address diverse challenges in temporal reasoning. Early research mainly relies on TimeML (Pustejovsky et al., 2003), focusing TimeX extraction and temporal relation extraction (Verhagen et al., 2007, 2010; UzZaman et al., 2013; Llorens et al., 2015; Miller et al., 2015; Mathur et al., 2021; Vashishtha et al., 2019). The advent of pre-trained language models (PLMs) has brought about commonsense reasoning as a tool to explore the world knowledge in models (Zhou et al., 2019; Qin et al., 2021; Dhingra et al., 2022). Recently, much attention has shifted towards event temporal reasoning (Chen et al., 2021; Tan et al., 2023; Zhu et al., 2023; Son and Oh, 2023; Chu et al., 2023b). Besides, Wang and Zhao (2023) introduces a unified form for assessing the temporal understanding capability of language models.
Distinguished from other datasets and benchmarks, TimeBench is multispectral, and closely aligned with intricate real-world scenarios, offering a comprehensive and hierarchical evaluation of LLM’s temporal reasoning abilities.
7.2 Large-scaled Language Model
In recent years, there has been rapid progress in the research of large-scale language models (LLM) (Zhao et al., 2023). They exhibit outstanding performance across a multitude of tasks without the need for fine-tuning (Brown et al., 2020; Kojima et al., 2022). Furthermore, they have achieved astonishing results in complex reasoning tasks, such as mathematical reasoning (Cobbe et al., 2021; Mishra et al., 2022) and logical reasoning (Yu et al., 2020; Liu et al., 2023). Moreover, some studies suggest that the chain-of-thought prompting can further enhance the model’s capabilities in complex reasoning scenario (Wei et al., 2022; Kojima et al., 2022; Chu et al., 2023a).
@@ 8 Conclusion
Time is a crucial dimension for understanding the world, and temporal reasoning inherently possesses complexity. The lack of a comprehensive benchmark makes it challenging to quantify LLMs’ temporal reasoning capabilities. In this work, we present TimeBench, a comprehensive and hierarchical benchmark for LLM temporal reasoning, tailored to mirror temporal reasoning in complex real-world scenarios. We conduct extensive experiments on state-of-the-art LLMs to investigate their temporal reasoning capabilities. Our findings indicate a substantial gap between state-of-the-art LLM and human performance, emphasizing the need for further research in this area. Moreover, we provide a meticulous analysis and discussion, outlining the current challenges that models face and suggesting potential directions for improvement.
Appendix A Appendix
A.1 Benchmark Details
We present the subtasks for each dataset in TimeBench, including data quantities, task formats, challenges faced by each task, and the capabilities required by models, as shown in Table 6.
A.2 Metrics for CTG
@@
Following SituatedGen (Zhang and Wan, 2023), we use BLEU-4 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr(Vedantam et al., 2015), and MATCH (Zhang and Wan, 2023) scores to metric the results of Constrained Text Generation. [^3] We utilize pycocoevalcap package to calucate BLEU-4, METEOR, ROUGE-L, CIDEr.
@@
The overall score is calculated as the sum of the above scores. We set the weight of CIDEr to $1/10$ for balance when summing.
@@
As the overall score $S$ does not represent a percentile, we proceeded to normalize the models’ scores to align with humans’ relative performance levels.
A.3 Full Results
The overall score is derived from the average of all corresponding metrics. For brevity, we omit some F1 scores in the main tables in the main text. Please refer to Table 7 for the full experimental results, and detailed metrics for SituatedGen can be found in Table 8.
A.4 Prompts
The prompt instructions are showcased in Figure 6. The demonstrations can be found from Figure 7 to Figure 15.
@@ A.4 Prompts
   - Order - 2,213 order
   - Duration - 2,332 duration
   - Conversion - 2,420 duration + time unit conversion
@@
MCTACO MCMAQ 852 Temporal Commonsense
TimeDial MCMAQ 1,446 Temporal Commonsense
DurationQA MCMAQ 687 Event Duration
@@
   - Explicit - 500 explicit, event-time reasoning
   - Implicit - 500 implicit, event-time reasoning
@@
   - Order - 400 event-time reasoning
   - Scope - 400 event-time reasoning
   - Counterfactual - 400 event-time reasoning
@@
   - l2 (e2t) - 839 event-time reasoning
   - l3 (e2e) - 1,037 event-event reasoning
@@
Table 6: The task formats and challenges in TimeBench.

### v2

@@ 7.1 Temporal Reasoning
There are numerous efforts addressing diverse challenges in temporal reasoning. Early research mainly relies on TimeML (Pustejovsky et al., 2003), focusing TimeX extraction and temporal relation extraction (Verhagen et al., 2007, 2010; UzZaman et al., 2013; Llorens et al., 2015; Miller et al., 2015; Mathur et al., 2021; Vashishtha et al., 2019). The advent of pre-trained language models (PLMs) has brought about commonsense reasoning as a tool to explore the world knowledge in models (Zhou et al., 2019; Qin et al., 2021; Dhingra et al., 2022). Recently, much attention has shifted towards event temporal reasoning (Chen et al., 2021; Tan et al., 2023; Wei et al., 2023). Han et al. (2021); Yang et al. (2023b); Son and Oh (2023); Chen et al. (2023) continuously pre-trains LLMs on time-aware data to elicit temporal reasoning, and Zhu et al. (2023); Su et al. (2023); Chu et al. (2023) explicitly represent temporal relationships using temporal graphs and timelines. Additionally, some works extend beyond text, evaluating temporal reasoning in structured tables and video domains (Gupta et al., 2023; Ko et al., 2023).
Some concurrent studies also analyze the temporal reasoning abilities of LLMs. Jain et al. (2023); Qiu et al. (2023) focus on temporal commonsense and Wang and Zhao (2023) introduces a unified form for accessing the overall abilities.
Distinguished from other works, TimeBench is multispectral, offering a comprehensive evaluation of LLM’s temporal reasoning abilities.
7.2 Large Language Models
In recent years, there has been rapid progress in the research of large language models (LLM) (Zhao et al., 2023). They exhibit outstanding performance across a multitude of tasks without the need for fine-tuning (Brown et al., 2020; Kojima et al., 2022). Furthermore, they have achieved astonishing results in complex reasoning tasks, such as mathematical reasoning (Cobbe et al., 2021; Mishra et al., 2022) and logical reasoning (Yu et al., 2020; Liu et al., 2023). Moreover, some studies suggest that the chain-of-thought prompting can further enhance the model’s capabilities in complex reasoning scenarios (Wei et al., 2022; Kojima et al., 2022; Chu et al., 2024; Zhang et al., 2023).
@@ 8 Conclusion
Temporal reasoning entails inherent diversity and complexity. The lack of a comprehensive benchmark makes it challenging to quantify LLMs’ temporal reasoning capabilities. In this work, we present TimeBench, a comprehensive and hierarchical benchmark for LLM temporal reasoning, tailored to mirror temporal reasoning in complex scenarios. We conduct extensive experiments on state-of-the-art LLMs to investigate their temporal reasoning capabilities. Our findings indicate a substantial gap between state-of-the-art LLMs and human performance, emphasizing the need for further research in this area. Moreover, we provide a meticulous analysis and discussion, outlining the current challenges that models face and suggesting potential directions for improvement.
Limitations
TimeBench is a comprehensive benchmark to quantify the temporal reasoning capabilities of LLMs. While we have taken various factors into account, there are a few limitations. Firstly, our evaluation only applied prompt-based method under zero-shot and few-shot setting, lacking evaluations specifically tailored for models fine-tuned on the temporal domain. Secondly, the instructions and demonstrations were manually crafted, which may potentially lead to discrepancies in prompts interpretation among different LLMs. Thirdly, the dataset constituting the benchmark includes data from past years and a portion sourced from Wikipedia, which may contaminate the training corpus of LLMs.
Acknowledgements
The research in this article is supported by the National Key Research and Development Project (2021YFF0901602), the National Science Foundation of China (U22B2059, 62276083), and Shenzhen Foundational Research Funding (JCYJ20200109113441941), Major Key Project of PCL (PCL2021A06). Ming Liu is the corresponding author.
Appendix A TimeBench Details
TimeBench features 3 major categories, 10 tasks and 15 subtasks, each with distinct challenges, totaling 19,000 instances. Detailed statistics are available in Figure 7 and Table 7.
A.1 Benchmark Construction
TimeX Arithmetic
(Tan et al., 2023) TimeX Arithmetic data is derived from the l1: time-time reasoning data in TempReason. We retain 4,000 instances, where time expressions are calculated with a minimum unit of one day.
TimeX NLI
(Thukral et al., 2021) The original data of TimeXNLI is in NLI format, including three sub-tasks, Temp-Order, Temp-Duration, and Cross-Unit Duration, including 6,140, 3,540, and 15,840 instances respectively. We conduct a random sampling of 2,213, 2,332 and 2,429 entries, resulting in a combined total of 6,965 instances.
MCTACO
(Zhou et al., 2019) The original MCTACO dataset consists of yes/no questions, containing 1,332 questions with 9,442 options. To guarantee that the questions are presented in a 4-way multi-select style, we initially remove questions that have less than four options. Subsequently, to ensure that each question has at least one correct option, we filter out questions where all options are labeled as "no". For each remaining question, we randomly sample four options, striving to maintain a balance between correct and incorrect options. In most cases, a question is accompanied by 2 correct and 2 incorrect options. A minority of questions have an option distribution of 1-3 or 3-1. After the aforementioned filtering process, we obtain 852 pieces of data in a 4-way multi-select format.
DurationQA
(Virgo et al., 2022) The original DurationQA has the same format as MCTACO, which consists of 694 questions with 4,868 options. Following the identical filtration procedure as MCTACO, we finally obtained a collection of 687 questions in a 4-way multi-select format.
TimeDial
(Qin et al., 2021) consists of 4-way multi-select instances in a two-person dialogue scenario. We leave the original data unaltered and simply randomize the sequence of options, yielding 1,446 pieces of 4-way multi-select instances.
SituateGen
(Zhang and Wan, 2023) SituatedGen includes 1,220 test cases, which span across two distinct reasoning domains: time and geography. We manually screen the original test data and retain those with clear time features for temporal reasoning evaluation, resulting in 115 instances.
TimeQA
(Chen et al., 2021) The original data of TimeQA includes two splits, Easy and Hard, with each question containing 20 Wikipedia paragraphs. The excessively long context may exceed the model’s maximum length limit and incur significant inference overhead. Therefore, we have reduced the context of the original data. For the paragraphs in the original data, we refer to those containing the answer as relevant paragraphs, and the rest as irrelevant paragraphs. For each question, we keep the first paragraph, all relevant paragraphs, and one random irrelevant paragraph as distractor. This ensures that most questions have at least three paragraphs. After that, we sample 500 pieces of data from those where the context length is less than 650 tokens. For both Easy and Hard splits, we apply the aforementioned filtration, resulting in 500 questions each, totaling 1,000 instances.
TempReason
(Tan et al., 2023) TempReason dataset contains 5,397 entries for l2 (event-time reasoning) and 4,426 entries for l3 (event-event reasoning). In the original dataset, each question corresponds with a text context and extracted facts. Similar to TimeQA, we apply a filter based on context length. We preserve questions with a context length between 300 and 600 tokens, yielding 839 and 1,037 instances, respectively. Notably, every remaining question is applicable to either context-based reasoning or fact-based reasoning.
MenatQA
(Wei et al., 2023) MenatQA consists of 999 data entries, formatted similarly to TimeQA, where each question is accompanied by several corresponding paragraphs. Following the paper’s proposed method, we modify the original data by incorporating the three time-sensitive factors: scope, order, and counterfactual. Subsequently, for each factor, we randomly sample 400 instances, resulting in a total of 1,200 data points.
TRACIE
(Zhou et al., 2021) The original TRACIE dataset consists of yes/no type questions, containing 4,248 test instances. We randomly sample 500 instances from the iid split in the test set.
A.2 Human Performance Evaluation
Unless otherwise stated, the results of human evaluation are derived from original dataset papers. Please refer to the corresponding paper for human evaluation details. TimeXNLI, Date Arith, and MCTACO are manually evaluated by three authors from the TimeBench team. Within each subtask, we randomly sample 50 instances, and the average of the performances by three human evaluators is considered the final human performance.
A.3 Task Formats
TimeBench is a multispectral benchmark, which features four different task formats.
Multi-Select Questions
Previous work utilizes the Multiple Choice (MC) form, which requires models to select the only correct answer from the options. However, this task form has shortcuts and may not truly reflect the model’s abilities. To address this, we employ the Multi-Select (M-S) task form, where the model needs to select all possible correct answers from the options provided. In our task, each question presents four options, with at most two of them being correct.
Natural Language Inference
is the task of determining the logical relationship between two pieces of text. Specifically, given a premise and a hypothesis, the model needs to determine whether the hypothesis can be inferred from the premise and output entailment, contradiction, or neutral. Our tasks focus on the entailment in temporal domains.
Free-form Reading Comprehension
requires models to answer questions based on the provided context, and the ground truth answer is free-form without pre-defined format restrictions.
Constrained Text Generation
refers to the task of generating text under certain constraints. The task is keyword-constrained text generation, where the model takes keywords as input and outputs sentences that include those keywords.
A.4 Evaluation Metrics
Accuracy is used for NLI and date arithmetic tasks. M-S tasks are evaluated using option-level EM and F1. FRC tasks (excluding date arithmetic) are assessed with token-level EM and F1. For CTG task, we take the average of multiple generation metrics, which are outlined as follows.
Metrics for SituatedGen
@@ A.4 Evaluation Metrics
Following SituatedGen (Zhang and Wan, 2023), we use BLEU-4 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2015), and MATCH (Zhang and Wan, 2023) scores to metric the results of CTG. [^3] We utilize pycocoevalcap package to calucate BLEU-4, METEOR, ROUGE-L, CIDEr.
@@
The overall score is calculated as the sum of the above scores. We set the weight of CIDEr to $1/10$ for balancing when summation.
@@
As the overall score $S$ does not represent a percentile, we proceed to normalize the models’ scores to align with humans’ relative performance levels.
Figure 7: The quantity and proportion of data for each task and its respective subtasks within TimeBench.
Appendix B Supplemental Materials
B.1 Models
ChatGPT-3.5/GPT-4
(Ouyang et al., 2022; OpenAI, 2023) ChatGPT is a chat model aligned through SFT and RLHF based on GPT-3 (Brown et al., 2020). GPT-4 is an upgraded version of ChatGPT with enhanced reasoning capabilities, making it the most powerful LLM. Unless otherwise stated, ChatGPT refers to gpt-3.5-turbo-0613 and GPT-4 refers to gpt-4-0613.
Llama2/Vicuna-1.5
(Touvron et al., 2023; Chiang et al., 2023) LLaMA2 is an open foundation model trained on 2T tokens with efficient grouped-query attention Ainslie et al. (2023). LLaMA2-chat is the official aligned model with SFT and RLHF, and Vicuna-1.5 is aligned with SFT only by the community [^4] https://lmsys.org/.
Baichuan2
(Yang et al., 2023a) is an open foundation model pre-trained on 2.6T tokens, which is competitive with LLaMA2. Baichuan2-chat is the official aligned model with SFT and RLHF.
Mistral
(Jiang et al., 2023) is a 7B open foundation model incorporating efficient grouped-query attention (Ainslie et al., 2023) and sliding windows attention (Beltagy et al., 2020). It achieves the strongest performance among models of its size, even surpassing LLaMA2-13B. Mistral-instruct is the officially aligned model with SFT only.
ChatGLM3
(Zeng et al., 2023) is an open-source bilingual LLM for Chinese and English, exhibiting competitive performance under 10B.
FLAN-T5
(Chung et al., 2022) is an open-source instruction model built on top of T5 (Raffel et al., 2020) through instruction fine-tuning.
B.2 Full Results
The overall score is derived from the average of all corresponding metrics. For brevity, we omit some F1 scores in tables in the main text. Please refer to Table LABEL:tab:full_results for the full experimental results. The full results of SituatedGen can be found in Table 8.
B.3 Prompts
The prompt formats are showcased in Figure 9. The demonstrations can be found from Figure 10 to 18.
Figure 8: Performance comparison between state-of-the-art LLMs. Up: GPT-4/3.5 and alignment models under zero-shot setting. Down: GPT-4/3.5 and base models under few-shot setting.
@@ B.3 Prompts
- Order - 2,213 order
- Duration - 2,332 duration
- Conversion - 2,420 duration + time unit conversion
@@
MCTACO M-S 852 Temporal Commonsense
TimeDial M-S 1,446 Temporal Commonsense
DurationQA M-S 687 Event Duration
@@
- Explicit - 500 explicit, event-time reasoning
- Implicit - 500 implicit, event-time reasoning
@@
- Order - 400 event-time reasoning
- Scope - 400 event-time reasoning
- Counterfactual - 400 event-time reasoning
@@
- l2 (e2t) - 839 event-time reasoning
- l3 (e2e) - 1,037 event-event reasoning
@@
Table 7: The statistics, task formats and challenges in TimeBench.
Method BLEU-4 METEOR ROUGE-L CIDEr MATCH Overall Norm
Human 39.9 40.4 56.3 397 98.1 274.4 100.0
GPT-4 8.23 31.27 28.84 38.45 90.41 162.59 59.25
+ FS 28.64 38.99 55.69 298.64 90.11 243.29 88.66
GPT-3.5 13.38 30.12 35.91 125.41 78.76 170.70 62.21
+ FS 27.24 33.77 51.18 282.75 76.54 217.01 79.08
LLaMA270b 5.15 13.62 15.83 22.07 31.79 68.60 25.00
+ FS 19.10 29.09 41.74 171.36 65.29 172.35 62.81
LLaMA213b 4.66 21.43 20.80 17.72 61.62 110.28 40.19
+ FS 15.15 27.49 37.55 138.13 64.94 158.93 57.92
LLaMA27b 2.77 13.46 14.69 14.34 34.83 67.18 24.48
+ FS 6.90 15.82 21.77 52.99 33.81 83.60 30.47
Baichuan213b 8.33 25.86 30.07 82.63 70.63 143.15 52.17
+ FS 15.79 30.23 40.96 169.14 71.01 174.91 63.74
Baichuan27b 5.17 21.99 23.73 44.80 59.85 115.22 41.99
+ FS 15.06 23.45 32.29 137.94 52.04 136.64 49.79
Vicuna1.513b 7.73 26.35 29.15 69.16 71.91 142.06 51.77
+ FS 6.85 18.66 25.99 92.96 46.19 106.99 38.99
Vicuna1.57b 6.29 24.34 26.91 46.90 68.84 131.07 47.77
+ FS 20.71 30.19 45.20 203.20 67.58 184.00 67.05
FLAN-T5 16.20 24.43 29.38 95.17 56.38 135.91 49.53
+ FS 12.88 30.38 36.27 92.20 76.44 165.19 60.20
Mistral7b 5.82 22.89 24.19 44.03 63.74 121.03 44.11
+ FS 18.96 29.02 43.15 185.61 63.24 172.93 63.02
ChatGLM36b 6.56 21.11 21.96 41.48 53.02 106.80 38.92
+ FS 10.53 24.17 33.44 124.50 56.94 137.53 50.12
[…]
Table 8: Full results of SituatedGen. Aligned models are under zero-shot setting by default. The top-3 results are bold. Methods with † are base models without alignment, under few-shot setting. We consider human performance as 100 points and normalize models’ results accordingly.